OpenStudy (anonymous):

Parallel and perpendicular lines.Write equation in slope intercept form for the line that passes through the given point and is parallel to the graph of each equation. (-2,-3) y=2x-1

OpenStudy (anonymous):

First off, parallel lines have the same slope. So you know that the slope of the line through that point is 2. Next, you need to use the point-slope form to eventually get to the slope-intercept form. The equation is y-y1=m(x-x1) Can you do it from there? Or is that where you have trouble?

OpenStudy (anonymous):

i kinda have trouble with it all, I had a sub today and it was he didn't explain it very well. I dont understand where i plug in the slope.

OpenStudy (anonymous):

how to get the slope*

OpenStudy (anonymous):

Oh okay. :) So an ordered pair is (x,y) right? So for x1 and y1, you plug in the ordered pair points. So it'd look like this: y-(-3)=2[x-(-3)] Then you'd simplify to look like y+3=2x+6 Then move the three to the right side of the plus sign to get the equation in y=mx+b form (m being slope, and b being the y-intercept of the line) Leaving y=2x+3 as your equation

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

Oh my bad, slope is expressed as 'm' So it's the number that is multiplied by the X Does that make sense?

OpenStudy (anonymous):

yea, so is that all?

OpenStudy (anonymous):

the parallel and perpendicular part confuses me

OpenStudy (anonymous):



OpenStudy (anonymous):

So parallel lines on the left, they aren't the same line, but they're headed in the same direction, so they have the same slope

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Perpendicular lines (on the right) intersect at a 90 degree angle, so the slopes of each are reciprocals. For example, if the slope of one line if 2/3, the perpendicular line's slope will be -3/2

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

So when you're given an equation and it says find the slope of a parallel line, you'd keep the same slope. For perpendicular, you'd flip the fraction and reverse the negative/positive sign Hope all that helps!

OpenStudy (anonymous):

thanks alot!

OpenStudy (anonymous):

No prob, lemme know if you have more questions :)